# Topological Data Analysis in smart manufacturing

### v2

@@ 7 Conclusion
For the survey, a transparent and rigorous methodology was employed to search for and identify relevant literature, ensuring the reproducibility of our findings. This approach yielded 35 works, which were analyzed. These studies were categorized into three distinct groups: Quality Control at the Product Level, Quality Control at the Process Level, and Manufacturing Engineering. Each study is reviewed, focusing on the data formats used and the specific TDA methods applied to each case.
@@ Acknowledgements
Martin Uray: Conceptualization, Methodology, Software, Validation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Barbara Giunti: Writing - Original Draft, Writing - Review & Editing, Visualization. Michael Kerber: Writing - Original Draft, Writing - Review & Editing. Stefan Huber: Writing - Original Draft, Writing - Review & Editing, Funding acquisition, Supervision.

### v3

@@ 7 Conclusion
For the survey, a transparent and rigorous methodology was employed to search for and identify relevant literature, ensuring the reproducibility of our findings. This approach yielded $34$ works, which were analyzed. These studies were categorized into three distinct groups: Quality Control at the Product Level, Quality Control at the Process Level, and Manufacturing Engineering. Each study is reviewed, focusing on the data formats used and the specific TDA methods applied to each case.
@@ Acknowledgements
Did I forget something? Did I do this correct? Martin Uray: Conceptualization, Methodology, Software, Validation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Barbara Giunti: Writing - Original Draft, Writing - Review & Editing, Visualization. Michael Kerber: Writing - Original Draft, Writing - Review & Editing. Stefan Huber: Writing - Original Draft, Writing - Review & Editing, Funding acquisition, Supervision.